# Quantitative Propagation of Chaos for 2D Viscous Vortex Model with General Circulations on the Whole Space

### v2

@@ 2.5. Solving the New ODE Hierarchy
The ODE hierarchy Proposition 2.1 is somehow an extension of the similar result in [55, Proposition 5], where the coefficients appearing in the ODE are either constants or exponential functions in time $t$, hence allowing S. Wang to apply the iterated integral estimates in [34]. However, in our cases, the growth in time function is more complicated from the logarithmic estimates, which forces us to deal with more general iterated integrals. The first half of the proof is similar to [55], but in the rest part we shall make a clearer computation.
Step 1: Change of Variables. Inspired by [55], we take the index
@@
$\displaystyle$== $\displaystyle\,\Big{(}\prod_{j=k}^{l-1}j\Big{)}\int_{0}^{t_{k}}\cdots\int_{0}^{t_{l-1}}\exp\Big{(}-\sum_{j=k}^{l-1}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}-l\varphi(t_{l})\Big{)}\Big{(}-kf(t_{k})f(t_{k+1})\cdots f(t_{l})\Big{)}$
@@
$\displaystyle$== $\displaystyle\,kC_{l-1}^{k}\int_{0}^{t}(1-e^{-\varphi(s)})^{l-k-1}e^{-\varphi(s)}f(s)\,\mathrm{d}s$
$\displaystyle$== $\displaystyle\,\frac{k}{l-k}C_{l-1}^{k}(1-e^{-\varphi(s)})^{l-k}\Big{|}_{0}^{t}$
$\displaystyle$== $\displaystyle\,C_{l-1}^{k-1}(1-e^{-\varphi(t)})^{l-k}.$
@@
$\displaystyle$==  $\displaystyle\,1-e^{-k\varphi(t)}-\sum_{j=k+1}^{l}C_{j-1}^{k-1}(1-e^{-\varphi(t)})^{j-k}e^{-k\varphi(t)}$
$\displaystyle$==  $\displaystyle\,1-e^{-k\varphi(t)}\sum_{j=0}^{l-k}C_{j+k-1}^{k-1}(1-e^{-\varphi(t)})^{j}.$
@@
$\displaystyle$== $\displaystyle\,\frac{l!}{(l-k)!(k-1)!}\int_{e^{-\varphi(t)}}^{1}z^{k-1}(1-z)^{l-k}\,\mathrm{d}z.$
@@
[…]
@@
$\displaystyle$== $\displaystyle\,\mathbb{P}(Y>e^{-\varphi(t)}),$
@@ 3. Regularity Estimates
[…]
@@
[…]
@@ 4.3. Logarithmic Hessian Estimate
[…]
@@
[…]
$\displaystyle$++ $\displaystyle\,\Big{(}\frac{M_{2}}{(1+t)^{2}}\phi+\psi^{\prime}+\frac{2M_{1}}{1+t}\psi+2\sigma\psi|\nabla\log\tilde{g}|^{2}+2\sigma C_{8}\log\tilde{g}+2\sigma C_{8}-\sigma C_{9}\Big{)}\frac{|\nabla\tilde{g}|^{2}}{\tilde{g}}.$

### v3

@@ 2.5. Solving the New ODE Hierarchy
The ODE hierarchy Proposition 2.1 is somehow an extension of the similar result in [55, Proposition 5], where the coefficients appearing in the ODE are either constants or exponential functions in time $t$, hence allowing S. Wang to apply the iterated integral estimates in [34]. However, in our cases, the growth in time function is more complicated from the logarithmic estimates, which forces us to deal with more general iterated integrals.
Step 1: Change of Variables. We take the index
@@
$\displaystyle=$ $\displaystyle\,\Big(\prod_{j=k}^{l-1}j\Big)\int_{0}^{t_{k}}\cdots\int_{0}^{t_{l-1}}\exp\Big(-\sum_{j=k}^{l-1}j\big(\varphi(t_{j})-\varphi(t_{j+1})\big)-l\varphi(t_{l})\Big)\Big(-kf(t_{k})f(t_{k+1})\cdots f(t_{l})\Big)$
@@
$\displaystyle=$ $\displaystyle\,kC_{l-1}^{k}\int_{0}^{t}(1-e^{-\varphi(s)})^{l-k-1}e^{-\varphi(s)}f(s)\,\mathrm{d}s$
$\displaystyle=$ $\displaystyle\,\frac{k}{l-k}C_{l-1}^{k}(1-e^{-\varphi(s)})^{l-k}\Big|_{0}^{t}$
$\displaystyle=$ $\displaystyle\,C_{l-1}^{k-1}(1-e^{-\varphi(t)})^{l-k}.$
@@
$\displaystyle=$  $\displaystyle\,1-e^{-k\varphi(t)}-\sum_{j=k+1}^{l}C_{j-1}^{k-1}(1-e^{-\varphi(t)})^{j-k}e^{-k\varphi(t)}$
$\displaystyle=$  $\displaystyle\,1-e^{-k\varphi(t)}\sum_{j=0}^{l-k}C_{j+k-1}^{k-1}(1-e^{-\varphi(t)})^{j}.$
@@
$\displaystyle=$ $\displaystyle\,\frac{l!}{(l-k)!(k-1)!}\int_{e^{-\varphi(t)}}^{1}z^{k-1}(1-z)^{l-k}\,\mathrm{d}z.$
@@
$\displaystyle\leq$ $\displaystyle\,\frac{1}{2}\sum_{l=k+1}^{\infty}l^{2}\Big(A_{k}^{l-1}(t)-A_{k}^{l}(t)\Big)$
$\displaystyle\leq$ $\displaystyle\,\frac{1}{2}\sum_{l=k+1}^{\infty}l^{2}B_{k}^{l}(t).$
@@
$\displaystyle=$ $\displaystyle\,\mathbb{P}(Y>e^{-\varphi(t)}),$
@@ 3. Regularity Estimates
$\displaystyle\leq\,$ $\displaystyle\frac{C}{t}\int_{{\mathbb{R}}^{2}}\exp\Big(-\frac{|x|^{2}}{8t+C_{0}}\Big)\exp\Big(-\frac{8t+C_{0}}{8tC_{0}}|y|^{2}\Big)\,\mathrm{d}y$
$\displaystyle\leq\,$ $\displaystyle\frac{C}{1+t}\exp\Big(-\frac{|x|^{2}}{8t+C_{0}}\Big).$
@@
$\displaystyle\geq$ $\displaystyle\,\frac{1}{C\lambda t}\int_{\mathbb{D}}\exp\Big(-C\frac{|x-y|^{2}}{t}-\lambda|y|^{2}\Big)\,\mathrm{d}y$
$\displaystyle\geq$ $\displaystyle\,\frac{1}{C\lambda t}\int_{\mathbb{D}}\exp\Big(-\frac{C\lambda}{C+\lambda t}|x|^{2}\Big)\exp\Big(-\frac{C+\lambda t}{t}|y|^{2}\Big)\,\mathrm{d}y$
$\displaystyle\geq$ $\displaystyle\,\frac{1}{C(1+t)}\exp\Big(-\frac{C\lambda}{C+\lambda t}|x|^{2}\Big).$
@@ 4.3. Logarithmic Hessian Estimate
$\displaystyle\leq$ $\displaystyle\,(M_{2}+2C_{5}\sigma+2C_{5}\sigma\log(1+M_{0})-C_{6}\sigma)\frac{|\nabla g|^{2}}{g}.$
@@
$\displaystyle\leq$ $\displaystyle\,\Big(\phi^{\prime}+\frac{4M_{1}+M_{2}}{1+t}\phi-\sigma\psi\Big)\frac{|\nabla^{2}\tilde{g}|^{2}}{\tilde{g}}$
$\displaystyle+$ $\displaystyle\,\Big(\frac{M_{2}}{(1+t)^{2}}\phi+\psi^{\prime}+\frac{2M_{1}}{1+t}\psi+2\sigma\psi|\nabla\log\tilde{g}|^{2}+2\sigma C_{8}\log\tilde{g}+2\sigma C_{8}-\sigma C_{9}\Big)\frac{|\nabla\tilde{g}|^{2}}{\tilde{g}}$.